# Hao Zheng Ma Jiang Chen Tian Chen Chen Xie

### 25 Mar 2026 v1

@@ Abstract
Regenerating singing voices with altered lyrics while preserving melody consistency remains challenging, as existing methods either offer limited controllability or require laborious manual alignment. We propose YingMusic-Singer, a fully diffusion-based model enabling melody-controllable singing voice synthesis with flexible lyric manipulation. The model takes three inputs: an optional timbre reference, a melody-providing singing clip, and modified lyrics, without manual alignment. Trained with curriculum learning and Group Relative Policy Optimization, YingMusic-Singer achieves stronger melody preservation and lyric adherence than Vevo2, the most comparable baseline supporting melody control without manual alignment. We also introduce LyricEditBench, the first benchmark for melody-preserving lyric modification evaluation. The code, weights, benchmark, and demos are publicly available at https://github.com/ASLP-lab/YingMusic-Singer.
@@ 1 Introduction
Singing Voice Synthesis (SVS) aims to generate human-like singing voices from musical scores, lyrics, and timbre references. Modern systems [1, 2, 3, 4, 5, 6] achieve high-fidelity synthesis, yet most rely on precisely annotated paired data associating each phoneme with an exact pitch contour and duration. While such fine-grained control is indispensable for professional music production, preparing these annotations creates a prohibitive barrier for an increasingly important use case, singing voice editing, which regenerates an existing singing voice with modified lyrics while preserving the original melodic and rhythmic structure. This capability is highly desirable for song adaptation, personalized cover generation, rapid prototyping of vocal arrangements, and multilingual song localization.
@@
Existing editing paradigms fall into two categories. The first adopts an in-context learning strategy that masks the region to be edited and regenerates it conditioned on the surrounding context and target lyrics [7, 8]. Although convenient, this approach is restricted to local segments and provides limited melody control. The second category, widely adopted in practice, relies on commercial SVS models such as Synthesizer V [^1] https://dreamtonics.com/synthesizerv and ACE Studio [^2] https://acestudio.ai, where users manually align modified lyrics with MIDI notes and durations before re-synthesizing the audio. This pipeline offers strong controllability but demands manual effort, and the complexity increases for tasks such as cross-lingual translation. In summary, existing approaches either depend on the surrounding context to recover melody or require manual alignment of word-level timestamps with melody information, limiting their flexibility and scalability. Several recent efforts address these challenges. Vevo2 [9] achieves melody-controllable generation, yet suffers from reduced intelligibility and poor melody adherence. SoulX-Singer [10] supports existing singing as melody input but still requires manual alignment of word-level timestamps, leaving the core burden unresolved. While these recent efforts narrow the gap, they either achieve suboptimal performance or still need manual alignment.
@@
To address these challenges, we propose YingMusic-Singer, a fully diffusion-based SVS model. First, YingMusic-Singer introduces a streamlined editing paradigm that synthesizes singing voices from only three inputs, namely an optional timbre reference clip, a singing clip providing the target melody, and the corresponding modified lyrics, without the need for manual alignment or precise annotation. Second, to mitigate limited phoneme generalization caused by the small scale and challenging vocal techniques in singing data, and to address the inherent trade-off between faithful lyric reproduction and melody adherence, we employ a curriculum training strategy combined with Group Relative Policy Optimization (GRPO) [11], enabling strong performance on both dimensions simultaneously. Third, we construct LyricEditBench based on GTSinger [12], the first benchmark for lyric modification evaluation under matched melody conditions, covering six common editing scenarios with balanced sampling for fair and comprehensive comparison. Experiments show that YingMusic-Singer outperforms Vevo2 [9], currently a comparable alignment-free alternative, in both melody preservation and lyric adherence. We will release model weights, inference code, and LyricEditBench to facilitate further research and advance the practical adoption of singing voice editing.
@@ 2.1 Architecture Overview
Figure 1: Overall architecture of YingMusic-Singer. Left: the training pipeline consisting of a Variational Autoencoder, a Melody Extractor, an IPA Tokenizer, and DiT-based conditional flow matching. Right: the GRPO training pipeline.
As shown in Figure 1, YingMusic-Singer generates singing voices at 44.1 kHz from three inputs: an optional timbre reference, a melody-providing singing clip, and corresponding modified lyrics. It comprises: (1) a Variational Autoencoder (VAE) following Stable Audio 2 [13], whose encoder $\mathcal{E}$ downsamples a stereo 44.1 kHz singing waveform $\mathbf{x}\in\mathbb{R}^{T\times 2}$ by a factor of 2048 into $\mathbf{z}=\mathcal{E}(\mathbf{x})\in\mathbb{R}^{T^{\prime}\times D}$, and whose decoder $\mathcal{D}$ reconstructs high-fidelity audio $\hat{\mathbf{x}}=\mathcal{D}(\mathbf{z})$ at inference; (2) a Melody Extractor built upon the encoder of a pretrained MIDI extraction model, whose intermediate representations naturally capture disentangled melody information. It produces $\mathbf{h}=\mathcal{M}(\mathbf{M})\in\mathbb{R}^{L\times D_{m}}$, which is then temporally interpolated to $\tilde{\mathbf{h}}\in\mathbb{R}^{T^{\prime}\times D_{m}}$ to match the VAE latent frame rate; (3) an IPA Tokenizer that converts both Chinese and English lyrics into a unified discrete phoneme sequence. To ensure the model correctly distinguishes between prompt and generation regions, avoiding phoneme omission or repetition at boundaries, we adopt sentence-level alignment following DiffRhythm [14]. Each lyric sentence is converted into an IPA subsequence and placed at its corresponding onset frame within a padded frame-level sequence of length $T^{\prime}$. The aligned sequence is passed through a learnable embedding layer to yield $\mathbf{e}\in\mathbb{R}^{T^{\prime}\times D_{e}}$. During inference, prompt lyrics are placed at the beginning of the sequence and target lyrics at the start of the masked region, requiring no timestamp annotation from the user; and (4) a DiT-based CFM backbone following F5-TTS [15].
@@ 2.2 Curriculum Training
To mitigate limited phoneme generalization caused by the small scale and challenging vocal techniques in singing data, YingMusic-Singer first undergoes TTS Pretraining without melody conditioning. The subsequent Singing Voice Supervised Fine-Tuning (SFT) stage has two phases. Phase 1 enables sentence-level alignment on singing data, allowing the model to adapt to the singing domain. Phase 2 activates melody conditioning and introduces a Centered Kernel Alignment (CKA) loss to enforce melody adherence. Given the predicted $v_{\theta}$ and melody $\tilde{\mathbf{h}}$, CKA measures their alignment via Gram matrices $\mathbf{K}=v_{\theta}v_{\theta}^{\top}$ and $\mathbf{L}=\tilde{\mathbf{h}}\tilde{\mathbf{h}}^{\top}$:
@@ 2.3 Group Relative Policy Optimization
While curriculum training achieves high performance, SFT Phase 2 simultaneously degrades PER, exposing a persistent trade-off. $\mathcal{L}_{\text{MSE}}$ targets holistic latent reconstruction and cannot isolate specific deficiencies for targeted optimization, while adjusting $\lambda$ in $\mathcal{L}_{\text{CKA}}$ only shifts the balance without resolving the trade-off itself. Moreover, inevitable noise in large-scale singing data, such as backing vocals and quality artifacts, caps model performance at the dataset ceiling. To overcome these limitations, reinforcement learning (RL) becomes essential. PPO requires a value network that is expensive to train. Offline methods such as DPO suffer from distribution shift, as pre-collected preference data becomes stale when the policy improves. GRPO [11] operates online and estimates baselines from within-group reward statistics, eliminating the value network while remaining efficient and stable.
Following recent efforts [16, 17, 18], we convert the deterministic ODE trajectory into an SDE but restrict stochastic steps to a bounded window, ensuring precise advantage attribution to exploratory steps. To prevent collapse toward a single reward dimension, we employ $M$ reward models jointly and compute the advantage for each sample as
@@
Melody Control Vevo2 [9] P $\downarrow$ 0.1378 0.1462 0.1545 0.1872 0.4409 0.4757 0.3352 0.3481 0.3812 0.3135 0.8019 0.5132
@@
Sing Edit Vevo2 [9] P $\downarrow$ 0.1290 0.1303 0.1596 0.1810 0.4111 0.4659 0.3414 0.3538 0.3531 0.2944 0.7680 0.4951
@@ 2.4 LyricEditBench
We build LyricEditBench from GTSinger [12] by removing all Paired Speech Group content, deduplicating audio via MD5 hashing, and excluding clips exceeding 15 seconds. DeepSeek V3.2 [19] then generates modified lyrics for each of the six modification types in Table 1. Given original lyrics and modification instructions, the LLM produces revised versions, with non-compliant outputs discarded, yielding 11,535 valid samples. Samples are classified by singer gender (male/female) and language (Chinese/English) into four categories, then organized by modification type. For each combination, we select 30 samples per singing technique (covering the six techniques in GTSinger) and 120 for the technique-free category, resulting in 300 per modification type per category and 7,200 test instances in total. For each instance, a timbre prompt of no more than 15 seconds is randomly drawn from the remaining audio pool, so each LyricEditBench instance comprises a melody reference clip, a timbre prompt, and the corresponding modified lyrics.
@@ 3 Experimental Setup
Dataset. The Chinese and English subsets of Emilia [20] are used for TTS pretraining. For Singing Voice SFT, internally licensed music tracks are processed by SongFormer [21] to segment structural boundaries and label function categories, discarding non-vocal segments. Vocal stems are then isolated using Mel-band RoFormer [22]. We retain clips between 2 and 30 seconds, splitting longer ones at sentence boundaries, ultimately obtaining 33,562.6 hours of singing data. The GRPO dataset is constructed by filtering SFT data with three criteria: ASR transcript verification, retaining only clips with a word error rate below 5%, speaker diarization via pyannote [23, 24], keeping only single-speaker clips, and a DNSMOS P808 quality score [25, 26] threshold of 3.5. This yields approximately 20,240 curated clips with balanced Chinese and English content. The test set in the proposed LyricEditBench is strictly excluded from training.
@@
Evaluation Metrics. We evaluate models with four objective and two subjective metrics. For objective evaluation, Phoneme Error Rate (PER) measures phoneme-level intelligibility, as singing exhibits lower semantic density, fewer contextual cues, and greater pronunciation variation than speech. Both Chinese and English clips are transcribed by singing-trained Qwen3-ASR-1.7B [27] and converted to phoneme sequences with tone markers removed. Speaker Similarity (SIM) follows F5-TTS [15], extracting speaker embeddings with a WavLM-large-based verification model and computing cosine similarity. F0 Pearson Correlation (F0-CORR) measures melody adherence via frame-wise Pearson correlation between F0 contours of generated and reference clips using RMVPE [28]. Vocal Score (VS) adopts VocalVerse2 [29] as a learned metric aligned with human perceptual preferences. For subjective evaluation, 120 samples uniformly sampled across task types and languages are rated by 30 listeners on two dimensions: Naturalness Mean Opinion Score (N-MOS) for overall perceptual quality and naturalness, and Melody Mean Opinion Score (M-MOS) for faithfulness to the reference melody, both on a 5-point scale.
@@
Implementation Details. We adopt the VAE from Stable Audio 2 [13] ($D=64$), the encoder of SOME [^3] https://github.com/openvpi/SOME as Melody Extractor ($D_{m}=128$, temporal dropout 0.1), and a DiT backbone following F5-TTS [15] (22 layers, 16 heads, hidden dim 1024, $D_{e}=512$). The full system has $\sim$727.3M parameters (453.6M CFM, 156.1M VAE, 117.6M Melody Extractor), trained on $\timesA8$00 80GB GPUs with DDP and bf16. Across all stages, 70%–100% of latent frames are randomly masked. TTS pretraining runs for 1M steps (batch duration 1.268h, lr 1e-4). Singing Voice SFT Phase 1 disables melody conditioning for 240K steps; Phase 2 enables it for 170K steps ($\lambda$ decayed from 0.3 to 0.01 over the first 2K steps; batch duration 1.69h, lr 2.5e-5). For GRPO, G=8G{=}8 candidates are scored by M=4M{=}4 equally weighted reward models (SDE noise a=0.8a{=}0.8, window $\mathcal{W}$ with $w_{\min}{=}1$, $w_{s}{=}8$, $\epsilon_{u}{=}0.01$, $\epsilon_{l}{=}0.002$, $\beta{=}1$), optimized for 1.2K steps (batch size 6, lr 7e-6) without CFG. Inference uses 32 ODE steps with CFG scale 3.

### 09 Apr 2026 v2

@@ Abstract
Regenerating singing voices with altered lyrics while preserving melody consistency remains challenging, as existing methods either offer limited controllability or require laborious manual alignment. We propose YingMusic-Singer-Plus, a fully diffusion-based model enabling melody-controllable singing voice synthesis with flexible lyric manipulation. The model takes three inputs: an optional timbre reference, a melody-providing singing clip, and modified lyrics, without manual alignment. Trained with curriculum learning and Group Relative Policy Optimization, YingMusic-Singer-Plus achieves stronger melody preservation and lyric adherence than Vevo2, the most comparable baseline supporting melody control without manual alignment. We also introduce LyricEditBench, the first benchmark for melody-preserving lyric modification evaluation. The code, weights, benchmark, and demos are publicly available at https://github.com/ASLP-lab/YingMusic-Singer-Plus.
@@ 1 Introduction
Singing Voice Synthesis (SVS) aims to generate human-like singing voices from musical scores, lyrics, and timbre references. Modern systems [DBLP:conf/aaai/Liu00CZ22, DBLP:conf/acl/HeLYHCLZ23, DBLP:conf/aaai/ZhangHLHXCDHZ24, DBLP:conf/iscslp/WangBDXGL24, DBLP:conf/slt/YuSWTW24, DBLP:conf/acl/ZhangGPYZ0W0025] achieve high-fidelity synthesis, yet most rely on precisely annotated paired data associating each phoneme with an exact pitch contour and duration. While such fine-grained control is indispensable for professional music production, preparing these annotations creates a prohibitive barrier for an increasingly important use case, singing voice editing, which regenerates an existing singing voice with modified lyrics while preserving the original melodic and rhythmic structure. This capability is highly desirable for song adaptation, personalized cover generation, rapid prototyping of vocal arrangements, and multilingual song localization.
@@
Existing editing paradigms fall into two categories. The first adopts an in-context learning strategy that masks the region to be edited and regenerates it conditioned on the surrounding context and target lyrics [lei2024songcreator, yang2025songeditor]. Although convenient, this approach is restricted to local segments and provides limited melody control. The second category, widely adopted in practice, relies on commercial SVS models such as Synthesizer V [^1] https://dreamtonics.com/synthesizerv and ACE Studio [^2] https://acestudio.ai, where users manually align modified lyrics with MIDI notes and durations before re-synthesizing the audio. This pipeline offers strong controllability but demands manual effort, and the complexity increases for tasks such as cross-lingual translation. In summary, existing approaches either depend on the surrounding context to recover melody or require manual alignment of word-level timestamps with melody information, limiting their flexibility and scalability. Several recent efforts address these challenges. Vevo2 [zhang2025vevo2unifiedcontrollableframework] achieves melody-controllable generation, yet suffers from reduced intelligibility and poor melody adherence. SoulX-Singer [qian2026soulxsingerhighqualityzeroshotsinging] supports existing singing as melody input but still requires manual alignment of word-level timestamps, leaving the core burden unresolved. While these recent efforts narrow the gap, they either achieve suboptimal performance or still need manual alignment.
@@
To address these challenges, we propose YingMusic-Singer-Plus, a fully diffusion-based SVS model. First, YingMusic-Singer-Plus introduces a streamlined editing paradigm that synthesizes singing voices from only three inputs, namely an optional timbre reference clip, a singing clip providing the target melody, and the corresponding modified lyrics, without the need for manual alignment or precise annotation. Second, to mitigate limited phoneme generalization caused by the small scale and challenging vocal techniques in singing data, and to address the inherent trade-off between faithful lyric reproduction and melody adherence, we employ a curriculum training strategy combined with Group Relative Policy Optimization (GRPO) [guo2025deepseek], enabling strong performance on both dimensions simultaneously. Third, we construct LyricEditBench based on GTSinger [DBLP:conf/nips/ZhangPGLZWXLHWZ24], the first benchmark for lyric modification evaluation under matched melody conditions, covering six common editing scenarios with balanced sampling for fair and comprehensive comparison. Experiments show that YingMusic-Singer-Plus outperforms Vevo2 [zhang2025vevo2unifiedcontrollableframework], currently a comparable alignment-free alternative, in both melody preservation and lyric adherence. We will release model weights, inference code, and LyricEditBench to facilitate further research and advance the practical adoption of singing voice editing.
@@ 2.1 Architecture Overview
Figure 1: Overall architecture of YingMusic-Singer-Plus. Left: the training pipeline consisting of a Variational Autoencoder, a Melody Extractor, an IPA Tokenizer, and DiT-based conditional flow matching. Right: the GRPO training pipeline.
As shown in Figure 1, YingMusic-Singer-Plus generates singing voices at 44.1 kHz from three inputs: an optional timbre reference, a melody-providing singing clip, and corresponding modified lyrics. It comprises: (1) a Variational Autoencoder (VAE) following Stable Audio 2 [DBLP:conf/ismir/EvansPCZTP24], whose encoder $\mathcal{E}$ downsamples a stereo 44.1 kHz singing waveform $\mathbf{x}\in\mathbb{R}^{T\times 2}$ by a factor of 2048 into $\mathbf{z}=\mathcal{E}(\mathbf{x})\in\mathbb{R}^{T^{\prime}\times D}$, and whose decoder $\mathcal{D}$ reconstructs high-fidelity audio $\hat{\mathbf{x}}=\mathcal{D}(\mathbf{z})$ at inference; (2) a Melody Extractor built upon the encoder of a pretrained MIDI extraction model, whose intermediate representations naturally capture disentangled melody information. It produces $\mathbf{h}=\mathcal{M}(\mathbf{M})\in\mathbb{R}^{L\times D_{m}}$, which is then temporally interpolated to $\tilde{\mathbf{h}}\in\mathbb{R}^{T^{\prime}\times D_{m}}$ to match the VAE latent frame rate; (3) an IPA Tokenizer that converts both Chinese and English lyrics into a unified discrete phoneme sequence. To ensure the model correctly distinguishes between prompt and generation regions, avoiding phoneme omission or repetition at boundaries, we adopt sentence-level alignment following DiffRhythm [ning2025diffrhythmblazinglyfastembarrassingly]. Each lyric sentence is converted into an IPA subsequence and placed at its corresponding onset frame within a padded frame-level sequence of length $T^{\prime}$. The aligned sequence is passed through a learnable embedding layer to yield $\mathbf{e}\in\mathbb{R}^{T^{\prime}\times D_{e}}$. During inference, prompt lyrics are placed at the beginning of the sequence and target lyrics at the start of the masked region, requiring no timestamp annotation from the user; and (4) a DiT-based CFM backbone following F5-TTS [DBLP:conf/acl/ChenN0DWZ0025].
@@ 2.2 Curriculum Training
To mitigate limited phoneme generalization caused by the small scale and challenging vocal techniques in singing data, YingMusic-Singer-Plus first undergoes TTS Pretraining without melody conditioning. The subsequent Singing Voice Supervised Fine-Tuning (SFT) stage has two phases. Phase 1 enables sentence-level alignment on singing data, allowing the model to adapt to the singing domain. Phase 2 activates melody conditioning and introduces a Centered Kernel Alignment (CKA) loss to enforce melody adherence. Given the predicted $v_{\theta}$ and melody $\tilde{\mathbf{h}}$, CKA measures their alignment via Gram matrices $\mathbf{K}=v_{\theta}v_{\theta}^{\top}$ and $\mathbf{L}=\tilde{\mathbf{h}}\tilde{\mathbf{h}}^{\top}$:
@@ 2.3 Group Relative Policy Optimization
While curriculum training achieves high performance, SFT Phase 2 simultaneously degrades PER, exposing a persistent trade-off. $\mathcal{L}_{\text{MSE}}$ targets holistic latent reconstruction and cannot isolate specific deficiencies for targeted optimization, while adjusting $\lambda$ in $\mathcal{L}_{\text{CKA}}$ only shifts the balance without resolving the trade-off itself. Moreover, inevitable noise in large-scale singing data, such as backing vocals and quality artifacts, caps model performance at the dataset ceiling. To overcome these limitations, reinforcement learning (RL) becomes essential. PPO requires a value network that is expensive to train. Offline methods such as DPO suffer from distribution shift, as pre-collected preference data becomes stale when the policy improves. GRPO [guo2025deepseek] operates online and estimates baselines from within-group reward statistics, eliminating the value network while remaining efficient and stable.
Following recent efforts [DBLP:journals/corr/abs-2505-05470, DBLP:journals/corr/abs-2507-21802, wang2026flowsegrpotrainingflowmatching], we convert the deterministic ODE trajectory into an SDE but restrict stochastic steps to a bounded window, ensuring precise advantage attribution to exploratory steps. To prevent collapse toward a single reward dimension, we employ $M$ reward models jointly and compute the advantage for each sample as
@@
Melody Control Vevo2 [zhang2025vevo2unifiedcontrollableframework] P $\downarrow$ 0.1378 0.1462 0.1545 0.1872 0.4409 0.4757 0.3352 0.3481 0.3812 0.3135 0.8019 0.5132
@@
Sing Edit Vevo2 [zhang2025vevo2unifiedcontrollableframework] P $\downarrow$ 0.1290 0.1303 0.1596 0.1810 0.4111 0.4659 0.3414 0.3538 0.3531 0.2944 0.7680 0.4951
@@ 2.4 LyricEditBench
We build LyricEditBench from GTSinger [DBLP:conf/nips/ZhangPGLZWXLHWZ24] by removing all Paired Speech Group content, deduplicating audio via MD5 hashing, and excluding clips exceeding 15 seconds. DeepSeek V3.2 [DBLP:journals/corr/abs-2512-02556] then generates modified lyrics for each of the six modification types in Table 1. Given original lyrics and modification instructions, the LLM produces revised versions, with non-compliant outputs discarded, yielding 11,535 valid samples. Samples are classified by singer gender (male/female) and language (Chinese/English) into four categories, then organized by modification type. For each combination, we select 30 samples per singing technique (covering the six techniques in GTSinger) and 120 for the technique-free category, resulting in 300 per modification type per category and 7,200 test instances in total. For each instance, a timbre prompt of no more than 15 seconds is randomly drawn from the remaining audio pool, so each LyricEditBench instance comprises a melody reference clip, a timbre prompt, and the corresponding modified lyrics.
@@ 3 Experimental Setup
Dataset. The Chinese and English subsets of Emilia [DBLP:conf/slt/HeSWLGHLYLSWCZW24] are used for TTS pretraining. For Singing Voice SFT, internally licensed music tracks are processed by SongFormer [hao2025songformerscalingmusicstructure] to segment structural boundaries and label function categories, discarding non-vocal segments. Vocal stems are then isolated using Mel-band RoFormer [DBLP:journals/corr/abs-2310-01809]. We retain clips between 2 and 30 seconds, splitting longer ones at sentence boundaries, ultimately obtaining 33,562.6 hours of singing data. The GRPO dataset is constructed by filtering SFT data with three criteria: ASR transcript verification, retaining only clips with a word error rate below 5%, speaker diarization via pyannote [DBLP:conf/interspeech/PlaquetB23, DBLP:conf/interspeech/Bredin23], keeping only single-speaker clips, and a DNSMOS P808 quality score [reddy2021dnsmos, reddy2022dnsmos] threshold of 3.5. This yields approximately 20,240 curated clips with balanced Chinese and English content. The test set in the proposed LyricEditBench is strictly excluded from training.
@@
Evaluation Metrics. We evaluate models with four objective and two subjective metrics. For objective evaluation, Phoneme Error Rate (PER) measures phoneme-level intelligibility, as singing exhibits lower semantic density, fewer contextual cues, and greater pronunciation variation than speech. Both Chinese and English clips are transcribed by singing-trained Qwen3-ASR-1.7B [shi2026qwen3asrtechnicalreport] and converted to phoneme sequences with tone markers removed. Speaker Similarity (SIM) follows F5-TTS [DBLP:conf/acl/ChenN0DWZ0025], extracting speaker embeddings with a WavLM-large-based verification model and computing cosine similarity. F0 Pearson Correlation (F0-CORR) measures melody adherence via frame-wise Pearson correlation between F0 contours of generated and reference clips using RMVPE [DBLP:conf/interspeech/WeiCDC23]. Vocal Score (VS) adopts VocalVerse2 [DBLP:conf/mm/WangYGLQLCFDZ25] as a learned metric aligned with human perceptual preferences. For subjective evaluation, 120 samples uniformly sampled across task types and languages are rated by 30 listeners on two dimensions: Naturalness Mean Opinion Score (N-MOS) for overall perceptual quality and naturalness, and Melody Mean Opinion Score (M-MOS) for faithfulness to the reference melody, both on a 5-point scale.
@@
Implementation Details. We adopt the VAE from Stable Audio 2 [DBLP:conf/ismir/EvansPCZTP24] ($D=64$), the encoder of SOME [^3] https://github.com/openvpi/SOME as Melody Extractor ($D_{m}=128$, temporal dropout 0.1), and a DiT backbone following F5-TTS [DBLP:conf/acl/ChenN0DWZ0025] (22 layers, 16 heads, hidden dim 1024, $D_{e}=512$). The full system has $\sim$727.3M parameters (453.6M CFM, 156.1M VAE, 117.6M Melody Extractor), trained on $\timesA8$00 80GB GPUs with DDP and bf16. Across all stages, 70%–100% of latent frames are randomly masked. TTS pretraining runs for 1M steps (batch duration 1.268h, lr 1e-4). Singing Voice SFT Phase 1 disables melody conditioning for 240K steps; Phase 2 enables it for 170K steps ($\lambda$ decayed from 0.3 to 0.01 over the first 2K steps; batch duration 1.69h, lr 2.5e-5). For GRPO, G=8G{=}8 candidates are scored by M=4M{=}4 equally weighted reward models (SDE noise a=0.8a{=}0.8, window $\mathcal{W}$ with $w_{\min}{=}1$, $w_{s}{=}8$, $\epsilon_{u}{=}0.01$, $\epsilon_{l}{=}0.002$, $\beta{=}1$), optimized for 1.2K steps (batch size 6, lr 7e-6) without CFG. Inference uses 32 ODE steps with CFG scale 3.